How do I find the opposite sine of -pi divided by 3 without using a calculator or the opposite sine of the positive sine of 2pi divided by 3?
Answer by stanbon(75887)

How do I find the opposite sine of -pi divided by 3 without using a calculator or the opposite sine of the positive sine of 2pi divided by 3?
----------
find the opposite sine of -pi divided by 3 without using a calculator
sin(-pi/3) = -sin(pi/3) = -sin(60 degrees) = -1/sqrt(3)
----
sin((2/3)pi) = sin(120 degrees) = sin(30) = 1/2
---------------------
Cheers,
Stan H.
=====================